OpenStudy (anonymous):

6^5x=20 round to the nearest-thousandth

OpenStudy (anonymous):

0.00257 = 0.0026 = 0.003. 0.003 is the answer :-)

OpenStudy (anonymous):

thanks

OpenStudy (triciaal):

6^5x =20 5xlog6 = log 20 (log 20/log 6)/5 = 1.301/.778 = 1.6722/5 = 0.334

OpenStudy (anonymous):

thanks

OpenStudy (triciaal):

WELCOME